### v1

# Squish Jamming

Abstract
A wide range of disordered materials, from biological to geological assemblies, feature discrete elements undergoing large shape changes. How significant geometrical variations at the microscopic scale affect the response of the assembly, in particular rigidity transitions, is an ongoing challenge in soft matter physics. However, the lack of a model granular-like experimental system featuring large and versatile particle deformability impedes advances. Here, we explore the oscillatory shear response of a sponge-like granular assembly composed of highly compressible elastic rings. We highlight a progressive rigidity transition, switching from a fluid-like to a solid-like response by increasing density or decreasing shear amplitude. The rearranging fluid state consists of crystal clusters separated by melted regions; in contrast, the solid state remains amorphous and absorbs all imposed shear elastically. We rationalise this transition by uncovering an effective, attractive shear force between rings that emerges from a friction-geometry interplay. If friction is sufficiently high compared to shear, the extent of the contacts between rings, captured analytically by elementary geometry, controls the rigidity transition.
Particulate assemblies, such as granular media, can switch their mechanical behaviour between a solid and a fluid-like response depending on their density, applied stress or interactions. This solid-fluid transition is of tremendous importance across length scales and disciplines, with the elemental particles varying in size, shape, and stiffness. Such jamming or yielding transitions are, for instance, underlying the onset of geological flows [1], the transport and clogging in industrial food processing [2], and are at the heart of numerous biological processes [3].
For rigid particles without internal degrees of freedom, the rigidity transition is mainly controlled by the average number of constraints imposed by the number of neighbours [4], including for non-spherical particles [5]. Even when particles are softer and can moderately deform their shape upon contact, such as emulsions or foams, the number of contacts remains an essential factor [6, 7, 8]. However, particle deformations may reach up to the order of the particle size for a broader class of materials, from biological tissues to porous soils. By adapting their shape to external constraints, such squishy particles [9] acquire additional degrees of freedom, potentially affecting the fluid-solid transition ingredients.
Recent works highlight the role of shape and deformability in the jamming transition of highly deformable particles, both in thermal [10, 11] and athermal [12, 13, 14] assemblies. Dedicated tools have also been developed to quantify the large shape changes that cells undergo while embedded in a deforming tissue [15]. The shape of the cells in epithelial tissues is one of the essential parameters governing a rigidity transition [16, 17, 18]. In geological flows, the shape of constitutive particles undergoes significant changes primarily via breaking [19], which in turn have a critical effect on the rheology [20] and jamming transition [21, 1]. With very soft grains that do not break, compaction far above jamming highlights that not only the coordination number increases with compaction, but also the size of the contacts [22, 23].
Despite these recent advances, a clear picture of how extreme deformability affects the fluid-solid transition is still lacking. It is an ongoing challenge in soft matter physics and related fields [24]. Although advanced numerical methods, such as the vertex model [18], the material point method [25] and phase field approach [26, 27], have been used, numerical investigations remain incredibly challenging owing to the non-linear couplings between geometry and interactions at the local scale, and a large number of elements. As such, a simple experimental model featuring large deformation at the grain scale would be essential to explore this question. One may leverage the recent advances in granular metamaterials [28] and metafluids [29], whose properties are governed by geometrical non-linearities at the particle scale, to control grains compressibility.
Here, we introduce assemblies of compressible elastic rings as an experimental model to explore the role of large grain deformations in the rheology of disordered media. Via oscillatory shear experiments, we highlight a solid-fluid transition controlled by the shear amplitude and the density. While the transition is characterised by the absence or presence of structural rearrangements, compression-induced large geometrical deformations help the assembly absorb shear without rearranging in the solid phase. Finally, we argue that this transition is mainly driven by the size of the frictional contacts between the compressed rings via effective adhesive interactions under shear, which stem from the interplay between geometrical deformations and friction.
I Fluid-solid transition in a sheared assembly of compressible rings
Figure 1: Fluid-solid transition under oscillatory shear of an assembly of compressible rings. a, Schematic of the experimental setup. A ring assembly is confined within a 2D cell with a fixed area. Oscillatory shear $\gamma=(\gamma_{0}/2)\sin\left(\dot{2\gamma_{0}}t/\gamma_{0}\right)$ is applied to the assembly via lateral motion of the top arm while the bottom one is fixed. Both top and bottom arms are decorated with half-solid disks of the same dimension as the rings to transmit shear. A transparent acrylic plate placed on top keeps the rings in the plane. b, Experimental and simulated (FEM) indentation response $f_{y}(\delta_{y})$, and c, associated stiffness $\frac{\mathrm{d}f_{y}}{\mathrm{d}\delta_{y}}$ of a single ring compressed between two flat plates. The experimental curve is the averaged response over five different rings from the assembly; upon loading and unloading, the shaded area shows the minimal and maximal values reached by the rings during the compression cycle. The corresponding FEM response is made via a 3D simulation of the neo-Hookean material model with Young’s modulus $E=1\,\mathrm{MPa}$ and Poisson’s ratio $\nu=0.45$, reproducing the ring and indentation geometry. The experiments and FEM, which agree quantitatively, feature a slight softening upon indentation, leading to solid foam-like deformation heterogeneities upon compression of the ring assembly. d, Top view of the ring assembly ($\Phi=1.04$ and $\gamma_{0}=0.53$) and representative of the pictures used for image analysis. Here, each tracked ring is coloured over its detected area by the value of its irreversible displacement $\log(\Delta u/r)$. The corresponding displacement is shown by a red arrow (up-scaled by a factor $4$). e, For each ring, we measure via image analysis its position (geometric centre), shape, neighbours, and extent of each contact (length of the blue segments). We define two neighbourhood types: Voronoi neighbours share a common edge in the Voronoi tessellation, and touching neighbours are in contact with a finite contact extent $\ell_{c}$. The thin orange silicone layer is applied in the inner part of the ring to help with detection and further imaging processes. See also Supplementary Video 2. f, Shear response $F_{X}(\gamma)$ of the assembly for $\Phi=1.00$ and for shear amplitude $\gamma_{0}$ ranging from $0.07$ to $0.7$. Six cycles in the steady state are displayed for each $\gamma_{0}$. g, Intensity of the averaged irreversible displacement $\log(\left\langle{\Delta u}\right\rangle/r)$ for varying density $\Phi$ and shear amplitude $\gamma_{0}$. The average is calculated over all tracked rings and in the steady state. h, Map of the ratio between the loss and storage moduli of the assembly, computed from the shear response $F_{x}(\gamma)$ and averaged over the cycles in the steady state. The corresponding $F_{x}(\gamma)$ for $\Phi=1.00$ are shown in f, each coloured by its corresponding value of $\log(\langle G^{\prime\prime}\rangle/\langle G^{\prime}\rangle)$. In g,h, and subsequent maps, the white dotted line is a geometric prediction introduced thereafter (Eq. (1)), but now serves as a reference to help compare the maps and identify solid and fluid regions.
We consider a two-dimensional assembly of frictional, elastic, and ring-shaped grains (see Fig. 1a and Supplementary Video 1). The assembly is compressed into an imposed area and sheared by oscillatory displacement of the top boundary as shown in the setup schematic in Fig. 1a. The rings are all identical with a thickness $t=1.5\,\mathrm{mm}$, radius $r=3.3\,\mathrm{mm}$ and height $h=10\,\mathrm{mm}$. Rings are fabricated via moulding of polymerising liquid silicone (see Methods and Supplementary Fig. LABEL:S-fig_sup_ring), then coated with talc powder to avoid adhesion and ensure dry frictional interactions. The compression response of the rings to a loading-unloading cycle (Fig. 1b) attests to a perfect elasticity with no remaining permanent deformations and high reproducibility, and is well recovered by the finite element method (FEM) (see Methods for details). In addition, the onset of the compression response features a small stiffness decrease (Fig. 1c). Deformed rings being slightly softer, a compressed assembly of such rings develops heterogeneous deformations as shown in the top view displayed in Fig. 1d (see also Supplementary Video 1). This heterogeneous deformation upon compression is reminiscent of the mechanics of cellular materials [30, 31, 32], including solid foams [33], with the critical difference that here the unit cells do not adhere each other and can therefore rearrange.
The assembly’s response to shear is explored by systematically varying the shear amplitude $\gamma_{0}$ and the density of rings $\Phi$. The oscillatory shear $\gamma$ is imposed at a fixed maximum shear rate of $\dot{\gamma_{0}}=1.33\times 10^{-2}\,\mathrm{s^{-1}}$, low enough to ensure a quasi-static regime where inertial or vibrational effects are negligible. The density is adjusted by varying the number of rings in the shear cell and quantified by $\Phi=N\pi r_{h}^{2}/A$ with $r_{h}=r+t/2$ the external radius of the rings, $N$ the number of rings and $A$ the area of the shear cell. This is the density that would correspond to the undeformed area of the rings, and we focus on the regime above the hexagonal packing density $\Phi_{h}=\frac{\pi}{2\sqrt{3}}\approx 0.91$. Initialisation is made by dispersing the rings in the cell area and then manually lowering and fixing the top arm at the desired height. The ring assembly is, therefore, initially in a mechanically quenched state. An acrylic plate on top of the assembly holds the rings in 2D, preventing out-of-plane buckling.
The mechanical response to shear is probed locally via image analysis and globally via a force sensor. Pictures from an overhanging camera are taken regularly during the cycles, such that the position, shape and neighbours of each ring are identified and tracked all along the shear cycles (Fig. 1e; see also Methods and Supplementary Video 2). A force sensor, placed between the driving screw and the top arm, records the force $F_{X}$ necessary to shear the assembly. The shear force $F_{X}$ for $\Phi=1.0$ and all investigated shear amplitudes is shown in Fig. 1f. Unless stated otherwise, the presented measures represent the system in the steady state (see Supplementary Table LABEL:S-suptab1 for the cycles used in the analysis). The high reproducibility of the shear response in Fig. 1f testifies that the assembly indeed reached a steady state.
We then assess whether the ring assembly responds more in a solid-like (jammed) or fluid-like (yielded) fashion. To assess this response, a classic measure for particulate systems under oscillatory loading is the irreversible displacement of the particles between two consecutive cycles [34, 35, 36, 37, 38, 39]. We define $\Delta u(i,c)$ as the norm of the irreversible displacement of ring $i$ occurring between the start and end of cycle $c$. In Fig. 1d, rings are coloured by their value of $\log\left(\Delta u/r\right)$ and the associated irreversible displacement is drawn with a red arrow (magnified by a factor 4). Even at such a high density, irreversible displacements occur but are highly heterogeneous, with values spanning orders of magnitude within the assembly. To explore how the irreversibility depends on the density and shear amplitude, for each parameter pair $(\Phi,\gamma_{0})$ considered, we compute $\langle\Delta u\rangle$, the averaged irreversible displacement over all tracked rings in the steady state. The resulting map is shown in Fig. 1g. Two zones can be identified: a fluid-like regime with significant irreversibility, which smoothly transitions to a solid-like zone with barely any irreversible displacement, by increasing density or decreasing shear (see also non-affine trajectories in Supplementary Video 3).
The transition between the fluid-like and solid-like features is observed not only in the rings’ local movement but also via the global mechanical response. From the shear response $F_{X}(\gamma)$ of the whole assembly, we compute for each cycle the first harmonic of the storage and loss moduli, $G^{\prime}$ and $G^{\prime\prime}$, respectively [40] (see separated map of $G^{\prime\prime}$ and $G^{\prime}$ in Extended Data Fig. 1). Identifying when the ratio $G^{\prime\prime}/G^{\prime}$ is greater than $1$ is a standard method to evaluate if the assembly yields or remains essentially solid [41]. The map of $\log(\left\langle{G^{\prime\prime}}\right\rangle/\left\langle{G^{\prime}}\right\rangle)$ shown in Fig. 1h, with $\langle{G^{\prime\prime}}\rangle,\langle{G^{\prime}}\rangle$ the moduli averaged over the cycles in the steady state, recovers qualitatively the solid- and fluid-like regions identified above from the irreversible displacements (Fig. 1g). Indeed, the region with high irreversible displacements coincides with a loss modulus significantly larger than the storage one and inversely for the dominantly reversible region.
Exploring the mechanical response to shear at very high density and shear amplitude, we identify a progressive jamming transition of the ring assembly by either reducing the shear amplitude or increasing the density. To understand the driving factors of this transition, we first quantify the structural properties of the assembly.
II A partially crystallised fluid and an amorphous solid
Figure 2: Crystallising fluid and amorphous solid. a, Left, probability density profiles of the hexagonal bond orientational order $\Psi_{6}$ for an assembly at $(\gamma_{0},\Phi)=(0.4,1.00)$ in the yielded region. The profiles are shown for the assembly before shear and after $20$, $40$ and $60$ shear cycles. Right, corresponding snapshots of the assembly after cycling, where each ring is coloured by its $\Psi_{6}$ value. The assembly shows a progressive growth of crystalline regions separated by melted ones. b, Same quantities as in a but for an assembly with higher density and same shear amplitude, $(\gamma_{0},\Phi)=(0.4,1.04)$, now located in the jammed region. Here, no significant structural changes are observed, and despite high shear, the assembly remains amorphous. See also Supplementary Video 4. c, Map of the proportion of rings belonging to crystal clusters (with $\Psi_{6}>0.75$) in the steady state. The corresponding snapshots of the assemblies in their steady states, with the rings coloured by their $\Psi_{6}$ values, are shown on the right for $\gamma_{0}=0.25$ and all investigated densities. See also Supplementary Video 5 for the time evolution of $\Psi_{6}$.
The presence or absence of irreversible displacements upon cyclic shear in the steady state highlights that ring rearrangements and associated structural deformations play a crucial role in the mechanical response of the assembly [42, 43]. We then measure the structural differences between the solid and fluid regions and how they evolve from their initial quenched state. With all rings having the same size, the configuration with the highest density and minimal shape deformation (and hence minimum elastic energy) is the hexagonal packing. We thus use the hexagonal bond orientational order $\Psi_{6}$ as the structural indicator of the assembly [44]. For each ring $j$, the hexagonal order is computed by $\Psi_{6}(j)=(1/n_{j})|{\sum_{k=1}^{n_{j}}\mathrm{e}^{6\mathrm{i}\alpha_{jk}}}|$ with $n_{j}$ the number of Voronoi neighbours and $\alpha_{jk}$ the angle between the line passing through the centre of the two rings and the $x$ axis (the direction of the shear). The value of $\Psi_{6}$ is $1$ for a ring in a perfect hexagonal lattice and continuously decreases to 00 as the structure deviates from the hexagonal one.
As the assembly accumulates shear cycles, we compare in Fig. 2 (see also Supplementary Video 4) the structural evolution of configurations for two cases located near the border between the jammed and yielded regions, with each case positioned on the opposite side of the border. After mechanical quenching, both configurations start in an amorphous state with a flat probability distribution $P(\Psi_{6})$, but their evolution with cycling strongly differ. In the yielded region, for $(\gamma_{0},\Phi)=(0.4,1.00)$ shown in Fig. 2a, cycling induces drastic structural changes with the progressive growth of distinct crystal clusters. The probability density increases close to $\Psi_{6}=1$ while it decreases for low values. In contrast, for larger densities within the jammed region (with $(\gamma_{0},\Phi)=(0.4,1.06)$ shown in Fig. 2b), no significant structural changes occur upon cycling, and the assembly remains in an amorphous phase. We then measure the crystallised portion of the assembly, or proportion of rings having $\Psi_{6}>0.75$, in the steady state over the parameter space (Fig. 2c). All the assemblies in the solid region feature almost no structural changes and remain amorphous, while the proportion of the crystal clusters decreases with increasing density in the fluid region (Fig. 2c; see also Supplementary Video 5). Rearranging rings within a given surface area tend to crystallise to reduce their overall elastic energy, as the hexagonal packing is the most compact structure without shape deformations. However, the growth of crystal clusters is limited by the presence of rings deformed by compression, which necessarily exist for high densities and prevent crystallisation.
The structure of the ring assembly in its steady state differs drastically between the yielded and jammed regions. The fluid-like region is characterised by hexagonal crystal clusters separated by a melted, disordered phase. This partial crystallisation is typical of granular media under sufficient shear where reorganisations lead to structural changes [45, 46, 47, 48]. Even if essentially crystallised, the yielded phase still displays a fluid-like mechanical response with irreversible displacements (Fig. 1g,h). In contrast, no crystallisation occurs in the jammed region, where the assembly remains amorphous almost without rearrangements. Even under high shear, the rings stick together to keep the same structure. In the following, we delve into the dynamics during shear cycles to investigate the interplay between structural and geometrical deformations of the rings.
III Large geometrical deformations absorb shear and hinder rearrangements
Figure 3: Competition between geometrical and structural deformations. a, The geometrical deformation of the rings is quantified by the shape index $q$. The evolution of $q$ within a cycle is shown for one ring in a $(\gamma_{0},\Phi)=(0.4,1.00)$ assembly. Each marker goes with an associated ring snapshot. We extract $\tilde{q}$ and $\Delta q$, the average and amplitude variation from the shape trajectory. b,c, Average shape of the assembly $\tilde{q}$ and shape variation $\Delta q$ as a function of shear amplitude for each density. The values are averaged over all tracked rings in the steady state. d,e, Map of the geometrical $S_{c}$ and topological $T_{c}$ strains of the assembly. The total strains here are the cumulative sums of the norm of the frame-to-frame strains estimated over the whole assembly and in the steady state.
We now focus on the deformation of rings during shear cycles to understand how yielded and jammed regions respond to shear. The geometry, of the rings is quantified by the shape index $q=(\mathrm{ring~{}perimeter})/\sqrt{\mathrm{ring~{}area}}$, a classic quantity used in highly deformable assemblies like epithelial tissues [16, 17] to assess how far a convex shape is from a circle. The shape index takes the value $q=2\sqrt{\pi}\approx 3.54$ for a perfect circle and increases when the ring gets deformed. As the rings are slender objects, they deform mostly by flexion rather than stretching. Thus, $q$ is mainly increased by reducing the area while the perimeter is almost unchanged. This contrasts with emulsions or tissues where most deformations originate from a change in perimeter and not the area [7, 27]. The perimeter and area of the rings are obtained via image analysis, and a Fourier decomposition is used to smooth the pixelated shape and reduce the shape to a few descriptors [27]. The resulting measure of $q$ during a shear cycle is shown in Fig. 3a for one ring within an assembly with parameters $(\gamma_{0},\Phi)=(0.4,1.00)$. For each cycle $c$ and tracked ring, we define two quantities from the shape trajectory: the average deformation $\tilde{q}=\left\langle{q}\right\rangle_{\mathrm{cycle}}$ and the geometrical fluctuation $\Delta q=\mathrm{max}(q)_{\mathrm{cycle}}-\mathrm{min}(q)_{\mathrm{cycle}}$.
In the steady state, the mean shape index and fluctuation are then averaged over all tracked rings, and their dependence on shear is compared for each density Fig. 3b,c. As expected, the average geometrical deformation $\tilde{q}$ increases steadily with the density Fig. 3b. For all densities, increasing shear amplitude drives significant shape fluctuations during shear cycles, and the higher the density, the more pronounced the geometrical changes (Fig. 3c). Thus, a higher density contributes, in addition to naturally increasing the average geometrical deformation, to promoting shape changes for a given imposed shear. Ring assemblies, therefore, use their compressibility and shape-changing abilities to absorb imposed external shear, and they do so more efficiently as the density increases.
To grasp the combined role of geometrical changes and structural rearrangements leading to jammed or yielded assemblies, we quantify these two aspects by dedicated strains derived for epithelial tissue deformations [15]. For each ring and between each time frame, we track the bond vectors between touching neighbours (Fig. 1e and Supplementary Video 6). The geometrical strain $S^{*}$ is derived from the change in bond length and orientation and quantifies changes in size and shape within the assembly. In contrast, the topological strain $T^{*}$, based on the appearance or creation of bonds due to rearrangements, captures the change in the topology of the contact network and then quantifies the structural deformations of the assembly. More details on these quantities are given in the Methods, and their step-by-step derivation is in the Appendix of Ref.[15]. The strains $S^{*}$ and $T^{*}$ are computed between each time frame and for the whole assembly. Their norm is then summed over all the cycles in the steady states to obtain the quantities $S_{c}=\int_{t}\norm{S^{*}}\mathrm{d}t$ and $T_{c}=\int_{t}\norm{T^{*}}\mathrm{d}t$ shown in Fig. 3d,e. From the map of $\log S_{c}$ over $(\gamma_{0},\Phi)$ in Fig. 3d, we observe that geometrical changes increase significantly with both shear amplitude and density as observed in Fig. 3c. By contrast, the map of $\log T_{c}$ in Fig. 3e shows that structural deformations, while increasing with shear amplitude, are diminished by higher densities. As a result, the region where changes in the topology of the contact network are significant coincides with the fluid region identified in Fig. 1g,h. Yet, the geometrical deformations vary regardless of the fluid or solid regions.
In contrast to confluent bi-dimensional active cell models where more significant deformations lead to yielding [18], here, deforming the assembly by increasing the density fosters jamming by absorbing shear and hindering structural rearrangements. We then investigate the ring-ring interactions to understand the mechanisms under which large shape deformations impede structural ones.
IV Geometry and friction concur to generate effective shear adhesion between rings.
Figure 4: Friction and geometrical deformations produce an effective shear adhesion between the rings. a, Ring-ring interaction experiments and FEM simulations. Two facing half-rings are compressed over a distance $\delta_{y}$ and then sheared laterally until they lose contact. Displayed are two FEM snapshots showing the contact pressure between the rings after compression, before (top) and after (bottom) a lateral shear displacement of $x=r$. See also Supplementary Video 7. b,c, The resulting shear interaction force $f_{x}$ is measured as a function of the lateral displacement $x$. Negative (positive) values indicate an attractive (repulsive) force between the rings. Experimental responses are averaged over $3$ distinct pairs, and shaded areas represent max and min values obtained from the different pairs. Panel b shows the shear response $f_{x}(x)$ for $\delta_{y}=950\,\mathrm{\mu m}$ and the friction coefficient $\mu=0.1,0.3,0.6$ varied in the FEM. The experimental response matches that of FEM for $\mu=0.6$. Panel c shows the shear response $f_{x}(x)$ for $\mu=0.6$, for increasing confinement $\delta_{y}=200,600,950\,\mathrm{\mu m}$. In this high friction regime, $f_{x}$ remains negative until the separation of the rings. We measure $x_{r}(\delta_{x},\mu)$, the displacement needed to switch from attractive ($f_{x}<0$) to repulsive ($f_{x}>0$) shear interactions. d, Experimental measure of the attractive-repulsive threshold shear $\gamma_{r}(\Phi)$, obtained by translating local quantities $(\delta_{y},x_{r})$ to global ones $(\Phi,\gamma_{r})$, together with results from FEM (dashed lines) and the geometrical predictions (dotted lines) for $\mu=0.1,0.3,0.6$. This curve $\gamma_{r}(\Phi)$ for $\mu=0.6$ is the one drawn in the phase diagrams shown so far (e.g., Fig. 1g,h), marking the fluid-solid transition of the assembly (fluid-like for $\gamma_{0}>\gamma_{r}$ and solid-like otherwise). e, $(\Phi,\gamma_{0})$ map of the difference between the imposed shear displacement $x_{0}$ and the measured maximal contact size $\langle{\ell_{c}^{\mathrm{max}}}\rangle$, averaged over all tracked rings in the steady state. The geometrical prediction and local measurements of the attractive-repulsive threshold $\gamma_{r}$ correspond to the region where $y_{0}\approx\langle{\ell_{c}^{\mathrm{max}}}\rangle$. In this high friction and large geometrical deformations limit, the extent of the contacts $\ell_{c}$ mainly controls the jamming transition.
We introduce a microscopic ersatz comprising two half-rings sequentially compressed and sheared to measure the interaction between rings. Then, we will link the sign of the shear force, which would promote or prevent structural rearrangements, to the assembly’s fluid-solid transition.
We measure, via FEM and experiments, the shear response of two clamped rings compressed by an amount $\delta_{y}$ (Fig. 4a and Supplementary Video 7). With FEM, we systematically vary the frictional interaction via the friction coefficient $\mu$ of the sheared interfaces (Fig. 4b) and the compression $\delta_{y}$ between the rings (Fig. 4c). Experimentally, we test rings taken from the assembly by applying the same compression values as in FEM. We find the best agreement between the experiments and the FEM for $\mu=0.6$, constituting a measure of the experimental friction coefficient. In the absence of friction, the shear force $f_{x}$ is instantly positive (repulsive) as elastic deformations push the rings away from each other. With friction, however, $f_{x}$ starts negative (attractive) and a finite lateral displacement $x_{r}(\delta_{y},\mu)$ is required to make it repulsive ($f_{x}>0$). The synergy of geometrical deformation and friction thus produces an attractive interaction, or effective adhesion, between the rings until sufficient shear is applied. The role of shape change to generate an effective adhesion under shear is reminiscent of the rheology of cohesive granular media [49], where the low stiffness of the grains also participates in producing a higher effective adhesion. We argue next that the presence or absence of this effective adhesion under shear is at the root of the assembly’s fluid or solid macroscopic response.
We now give a geometrical prediction of the finite lateral displacement $x_{r}$ necessary to switch the shear interactions force from attractive to repulsive, i.e., $f_{x}(x_{r})=0$. This distance is a function of the initial compression $\delta_{y}$ and the friction between the rings, $\mu$. Two cases emerge in the way $x_{r}(\delta_{y},\mu)$ is determined. In the first case, the rings remain in contact when $f_{x}$ switches from negative to positive (e.g., Fig. 4b for $\mu=0.1$ and $0.3$). In this case, assuming Amontons-Coulomb frictional properties and reducing the contacting interfaces to a point, one can directly link the sign of the shear force $f_{x}$ to the lateral displacement and obtain $x_{r}(\delta_{y},\mu)=\mu(2r_{h}-\delta_{y})$. By contrast, in the second case where $f_{x}$ remains negative until the rings are fully separated (e.g., Fig. 4c), $x_{r}$ is given by the extent of the initial overlap between the rings: $x_{r}(\delta_{y},\mu)=2r_{h}\sqrt{1-(1-\delta_{y}/(2r_{h}))^{2}}$. Now that we have a prediction for the microscopic ring-ring interactions, we must translate the compression $\delta_{y}$ and local lateral displacement $x$ to the global variable density $\Phi$ and shear $\gamma$. Assuming an hexagonal structure of density $\Phi_{h}$, an isotropic local overlap $\delta_{y}$ directly translates into a density by $\Phi(\delta_{y})=\Phi_{h}/(1-\delta_{y}/(2r_{h}))^{2}$. For the shear $\gamma$, too, by using its local counterpart, i.e., the local shear imposed between the two half rings, we obtain $\gamma=x/(2r_{h}-\delta_{y})$. Schematics of the underlying geometrical predictions are shown in Extended Data Fig. 2. By combining the expressions of $x_{r}(\delta_{y},\mu)$ with the expressions of $\Phi(\delta_{y})$ and $\gamma(y,\delta_{y})$, we can finally express the repulsive shear $\gamma_{r}$ at which the interactions become repulsive as a function of global variables $(\gamma,\Phi)$, as follows:
$\gamma_{r}(\Phi,\mu)=\begin{cases}\sqrt{\frac{\Phi}{\Phi_{h}}-1}&\mathrm{if}\,\,\gamma_{r}<\mu\\ \mu&\mathrm{if}\,\,\gamma_{r}\geq\mu\end{cases}$ (1)
In Fig. 4d, we compare $\gamma_{r}$ from the prediction against FEM and experimental measurements and find a perfect match. In the high friction case ($\gamma_{r}<\mu$), the threshold $\gamma_{r}(\Phi)$ is simply set by the extent of contacts, yielding $\gamma_{r}(\Phi)=\sqrt{\Phi/\Phi_{h}-1}$. Experimentally, the explored density range and the estimated friction coefficient $\mu\approx 0.6$ are such that we remain in this high friction limit. In soft spheres without shape changes, the jamming transition also features a scaling between yielding and density [50] but occurs at a much lower density and with a different scaling.
With repulsive interactions between rings, we expect structural rearrangements to be promoted and, inversely, hindered if interactions are attractive and act as an effective adhesive force. The shear $\gamma_{r}$ can then be viewed as a yield shear, where rearrangements leading to fluid-like response are expected for $\gamma_{0}>\gamma_{r}$, and a solid-like jammed assembly for $\gamma_{0}<\gamma_{r}$. Indeed, the line drawn by the geometric prediction of Eq. (1) qualitatively separates the fluid and solid-like regions in the $(\Phi,\gamma_{0})$ parameter space (Fig. 1g,h), which also accompany significant changes of the contact network topology (Fig. 3e). This geometric boundary $\gamma_{r}(\Phi)=\sqrt{\Phi/\Phi_{h}-1}$ is set by the extent of the contact between rings, which increases with density. To verify that the contact extent is a major factor controlling the fluid-solid transition, we measure ${\ell_{c}^{\mathrm{max}}}$, the largest contact length among all touching neighbours, for each ring and at each time step. We then compute the difference between the equivalent local lateral displacement amplitude $x_{0}=\gamma_{0}(2r_{h}-\delta_{y})=\gamma_{0}2r_{h}\sqrt{\Phi_{h}/\Phi}$ and $\langle{\ell_{c}^{\mathrm{max}}}\rangle$, averaged in the steady state (Fig. 4e). The geometrical prediction of $\gamma_{r}$ matches with the region where $x_{0}\approx\langle{\ell_{c}}^{\mathrm{max}}\rangle$. The fluid-solid regions are then defined by whether the shear is large enough to overcome the effective adhesion and move the rings by a distance greater than the maximal contact length. Therefore, in contrast with rigid granular media where the number of contacts controls the jamming transition [4], the transition with squishy grains in a high friction limit seems to be strongly controlled by the extent of the contacts (Extended Data Fig. 3).
V Discussion
We established that large shape deformation may radically modify the mechanisms by which an assembly yields. At high density, far into the jammed state, the number of contacts becomes irrelevant, and their extent, coupled with frictional interactions, becomes crucial. The contact size is intrinsically linked to the geometrical deformations of the elements, so the shape also has a crucial role in the rigidity transition. Geometry and friction concur to generate an attractive force between rings under shear, and sufficient shear must be applied to overcome this effective adhesion. In the high friction limit, the shear must be enough to separate the rings completely to overcome the adhesion; hence, a simple geometric formula (Eq. (1)) sets the transition between repulsive and attractive interactions. Below this transition (when the interactions are attractive), structural rearrangements are strongly hindered, leading to a jammed assembly.
However, these results are valid only within significant friction and shear limits. Our geometrical model implies that a frictionless assembly yields at an infinitesimal shear as $\gamma_{r}=\mu$. By considering only the interactions between two rings, we neglect collective effects, such as the caging [51], that would require a finite shear to yield even without friction. How caging is affected by significant shape changes is still an open question.
Another striking property of the ring assembly is the drastic structure change between the fluid and solid state. Counterintuitively, the fluid phase features highly crystallised portions, while the solid remains in an amorphous structure. Similarly to the link between active matter and sheared granular media [52], one may also draw a parallel between this fluid-like crystal under shear and active crystals [53, 54].
From an engineering perspective, assemblies of highly deformable elements combine frictional interactions and structural rearrangement to dissipate energy while keeping a significant elastic component [55]. Ring assemblies then have great potential as a tunable shock absorber, with the density serving as a knob to tune dissipation and response to a shock, keeping even a high dissipative potential in the solid and reversible state. To develop this axis, it would be essential to explore the 3D version of such assemblies and their response to inertial forces.
Acknowledgements.
Acknowledgments.
This work benefited from numerous enriching discussions with, but not limited to, A. Ikeda, N. Saito, S. Ishihara, Y. Lou, H. Ikeda, F. Lechenault, C. Scalliet, M. Henot, I. Cantat, B. Chakraborty, and B. Andreotti. We thank R. Tosaka for the construction of a prototype of the shear device. This work is supported in part by KAKENHI from the Japan Society for the Promotion of Science (Grant Nos.JP22KF0084, JP20H00128, JP19H05800, JP24K00593). S.P. acknowledges financial support from the Japanese Society for the Promotion of Science as a JSPS International Research Fellow.
Authors contribution.
S.P. and K.A.T. designed the experiments. S.P. performed the experiments, the numerical simulations, the data analysis, and conceived the model. All authors analysed and discussed the results and contributed to writing the paper.
Code and data availability
The Zenodo repository 10.5281/zenodo.10961344, contains the experimental data for the force and from image analysis for ring assemblies and single ring experiments. The raw images of the first five cycles for $\gamma_{0}=0.40$ and $\Phi=1.00$ and the MATLAB code to process them. It also contains the FEM data and associated COMSOL files.
Figure 1: Loss and storage moduli across the parameter space. The first harmonic of the loss and storage moduli are shown in a and b, respectively. The moduli are normalised by the Young modulus of the rings, $E=1\,\mathrm{MPa}$.
Figure 2: Schematic for the geometrical predictions. a, First case to estimate the lateral displacement $x_{r}$ which for which the lateral force $f_{x}$ becomes repulsive ($f_{x}>0$). The contact is reduced to a single point with an orientation $\alpha$ given by the line linking the two intersection points of the half-circles. Assuming frictional Coulomb interactions, the amplitude of tangential $\mathbf{T_{t}}$ and normal $\mathbf{T_{n}}$ contact forces are linked by the friction coefficient $\mu$. The lateral force $f_{x}$ is the resultant of the contact forces projected on the lateral axis: $f_{x}=(\mathbf{T_{n}}+\mathbf{T_{t}})\cdot\mathbf{e_{x}}$. The expression for $x_{r}$ is then found by solving the equation $f_{x}(x=x_{r})=0$. This prediction is valid if the two rings are still in contact at $x=x_{r}=\mu(2r_{r}-\delta_{y})$. b, If the rings are not in contact at $x=x_{r}=\mu(2r_{r}-\delta_{y})$, we have to consider the second case where $f_{x}$ remains attractive until complete separation of the rings. In that case, $x_{r}$ is the no-overlap distance given by the initial contact extent $\ell_{c}$ after compression. c, To link the global assembly density $\Phi$ and the local compression $\delta_{y}$, we assume an isotropic and homogeneous compression of an assembly initially in a hexagonal packing. The density is estimated by considering a central ring and the hexagon with corners at the centre of its $6$ neighbours and by taking the ratio between the area of the ring inside the hexagon ($1$ full ring + $6$ third of rings) and the area of the hexagon. In the hexagonal lattice, the hexagon edge is $2r_{h}$ and reduces to $2r_{h}-\delta_{y}$ under compression, while the area of the rings is constant ($\Phi$ is estimated from the area of undeformed rings). d, Shear is a scaleless quantity defined as the ratio between the lateral displacement and the height of the system. The shear defined for the assembly as $\gamma_{=}X/T$ is then identical to the one in the local interaction experiments, $\gamma=x/(2r_{h}-\delta_{y})$.
Figure 3: The average coordination number is irrelevant for highly compressed assemblies. a, Average of the coordination number $Z_{r}$ of the assemblies in the steady state. The coordination number of a ring is defined as the number of touching neighbours. No clear tendency with $\gamma_{0}$, $\Phi$, or the fluid/solid regions is noticeable. b, Reproduction of Fig. 4e, showing that the relative value between the shear distance $x_{0}$ and the averaged maximal contact length $\langle{\ell_{c}}^{\mathrm{max}}\rangle$ marks a clear separation between fluid and solid-like regions. c, local average lateral displacement $x_{0}$ imposed to the assembly over the parameter space $(\gamma_{0},\Phi)$. It is a deterministic quantity computed following $x_{0}=\gamma_{0}2r_{h}\sqrt{\Phi_{h}/\Phi}$. d, For each tracked ring in the steady state, the contact length $\ell_{c}$ with each of its touching neighbours is measured (see Methods). At each time-step, the highest $\ell_{c}$ of each ring is averaged over all rings, and then all time-steps to give $\langle{\ell_{c}}^{\mathrm{max}}\rangle$. While $x_{0}$ mainly increases with the shear $\gamma_{0}$, $\langle{\ell_{c}}^{\mathrm{max}}\rangle$ is increasing with density $\Phi$. The two values coincide around the fluid-solid separation line predicted by $\gamma_{r}(\Phi)=\sqrt{\Phi/\Phi_{h}-1}$.